# Exploring the ability of the Deep Ritz Method to model strain localization as a sharp discontinuity

### v1

@@ 7 Conclusions
We have used 1D and 2D numerical examples as proof of concept that plastic solids undergoing strain localization can be, in principle, modeled using the DRM combined with regularized strong discontinuity kinematics to approximate the displacement field. Furthermore, this result suggests that variational versions of PINNs hold promise for solving complex computational solid mechanics problems, such as materials with strain localization.
The present study sets the stage for subsequent developments on several fronts. The idea of describing strong discontinuities via trainable ANNs can be applied to many cases beyond softening plasticity, e.g., frictional-dilatant plasticity or damage, and with more general loading conditions. Moreover, while we have shown feasibility for the onset and evolution of fixed localization bands, applying the proposed scheme to general boundary value problems requires handling crack/band propagation. A possible way forward is to develop multiscale techniques with ANN-based discontinuities propagating through a macroscopic mesh.

### v2

@@ 7 Conclusions
We have used 1D and 2D numerical examples as proof of concept that plastic solids undergoing strain localization can be, in principle, modeled using energy minimization combined with regularized strong discontinuity kinematics to approximate the displacement field. In particular, we have been able to predict the onset and evolution of a localization band, including its position and the displacement jump magnitude, via trainable parameters in an ANN, addressing persistent challenges in finite element technologies endowed with sharp discontinuities. This result suggests that variational versions of PINNs hold promise for solving complex computational solid mechanics problems, such as materials with strain localization.
The present study sets the stage for subsequent developments on several fronts. Firstly, we emphasize that this study has not focused on developing a systematic and efficient numerical procedure but rather on proposing a potentially new paradigmatic approach to strain localization. Therefore, future work should focus on optimizing the computational cost, which is rather high for the training process, and streamlining the overall procedure. On the other hand, the idea of describing strong discontinuities via trainable ANNs can be applied to many cases beyond softening plasticity, e.g., frictional-dilatant plasticity or damage, and with more general loading conditions. Moreover, while we have shown feasibility for the onset and evolution of fixed localization bands, applying the proposed scheme to general boundary value problems requires handling crack/band propagation. A possible way forward is to develop multiscale techniques with ANN-based discontinuities propagating through a macroscopic mesh.
CRediT authorship contribution statement
Omar León: Methodology, Software, Formal analysis, Investigation, Writing - Original Draft, Writing - Review & Editing, Visualization. Víctor Rivera: Methodology, Software, Formal analysis, Investigation, Writing - Original Draft, Writing - Review & Editing, Visualization. Angel Vázquez-Patiño: Methodology, Software, Writing - Review & Editing, Visualization. Jacinto Ulloa: Validation, Formal analysis, Writing - Review & Editing. Esteban Samaniego: Conceptualization, Methodology, Validation, Formal analysis, Writing - Original Draft, Writing - Review & Editing, Supervision, Project administration.
Data availability
No data was used for the research described in the article.